In 1895, the first Putting Green Championship was held. The winner's prize money was 200. In 2014, the winner's check was 1,400,000. What was the percentage increase per year in the winner's check over this period? If the winner's prize increases at the same rate, what will it be in 2035
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1,400,000=200*(1+r/100)^119

(1,400,000/200)^(1/119)=(1+r/100)

(1+r/100)=1.077238207

r=(1.077238207-1)

=7.72%(Approx).

b.

A=$1,400,000*(1.077238207)^21

=$1,400,000*4.770311211

=$6,678,435.70(Approx).